# Pattern formation and spatiotemporal chaos in relativistic degenerate plasmas

### v1

@@ Abstract
We numerically study the nonlinear interactions of high-frequency circularly polarized electromagnetic (EM) waves and low-frequency electron-acoustic (EA) density perturbations driven by the EM wave ponderomotive force in relativistic plasmas with two groups of electrons–the population of relativistic degenerate dense electrons (bulk plasma) and the sparse relativistic nondegenerate (classical) electrons, and immobile singly charged positive ions. By pattern selection, we show that many solitary patterns can be generated and drenched through modulational instability of EM waves at different spatial length scales and that the EM wave radiation spectra emanating from compact astrophysical objects may not settle into stable envelope solitons but into different incoherent states, including the emergence of temporal and spatiotemporal chaos due to collisions and fusions among the patterns with strong EA wave emission. The appearance of these states is confirmed by analyzing the Lyapunov exponent spectra, correlation function, and mutual information. As a result, the redistribution of wave energy from initially excited many solitary patterns at large scales to a few new incoherent patterns with small wavelengths in the system occurs, leading to the onset of turbulence in astrophysical plasmas.
I Introduction
@@
The nonthermal radiation of intense electromagnetic (EM) waves (radio wave, $X$-rays, or $\gamma$-rays) can be observed from pulsating stars and during their collapse due to compression of their magnetospheres and a significant increase in the magnetic field. The electric field in this process accelerates charged particles, resulting in charged particle radiation when they move in the magnetic field [1]. Thus, strong high-frequency (hf) EM waves and high-density relativistic degenerate dense plasmas associated with compact objects (e.g., white dwarfs, neutron stars, active galactic nuclei, etc.) can be simultaneously produced. Such an intense EM wave could play a crucial role in characterizing the nature of the final radiation spectra that will emerge from these compact objects. In the latter, plasmas may consist of high-density relativistic degenerate electrons, contaminated by a small fraction of classical or non-degenerate electrons and positive ions forming only the neutralizing background plasma. They thus can support the propagation of low-frequency (lf) electron-acoustic waves (EAWs) [2, 3]. Since hf EM fields can scatter into lf EM waves on EAWs, the EM spectral range will change from the original radiation and carry essential information while interacting with EAWs for the self-modulation, soliton formation [2, 4], and the emergence of spatiotemporal chaos [5], leading to turbulence in astrophysical plasmas. Several authors have studied the nonlinear interactions of hf EM waves with electrostatic lf density perturbations for generating electromagnetic envelope solitons and wakefields [6, 7]. However, less effort has been paid to study the emergence of spatiotemporal chaos or EM wave turbulence due to these interactions in high-density regimes.
@@
Electromagnetic wave turbulence in astrophysical environments can involve interactions of several physical processes, including shocks, instabilities, and reconnection, and these interactions can result in some observable phenomena like plasma flares and jets, plasma eruptions, and sunspots. The EM wave turbulence can also be generated by other processes like linear conversion of wave modes on plasma density fluctuations, which involve resonant three-wave interactions, and fusion of electrostatic or EM waves. However, several attempts have also been made to understand wave turbulence in terms of nonlinear dynamics and chaos, which involve pattern formation and extensive and incoherent pattern dynamics in both space and time [8, 9]. We will follow this procedure in the present investigation. In this process, the most significant conservative version of wave turbulence is the Zakharov equations (See, e.g., [5]) that couple hf slowly varying electrostatic (e.g., Langmuir wave) or EM waves (linearly or circularly polarized) to slow electrostatic plasma density perturbations (e.g., electron-acoustic wave and ion-acoustic wave). Typically, decay processes (which may be an induced emission of some waves by others or nonlinear interactions of waves that result in the break-up of waves into others) accumulate wave energy into perturbations with long wavelengths. However, when this wave energy exceeds the modulational instability (MI) threshold, a transfer of energy or a redistribution of energy from initially excited large-scale modes to higher harmonic modes with small wavelengths may be possible. It is believed that some sort of chaotic process in a subsystem may be responsible for driving the energy transfer [5]. Nevertheless, even when the MI of wave envelopes is still present and the energy threshold is exceeded, the energy transfer may not occur, e.g., in the description of wave envelopes by a nonlinear Schröinger (NLS) equation. The latter is known to be an integrable version of the Zakharov equations in the adiabatic limit (very slow-time scales) in which case the MI gets saturated in the formation of envelope solitons [10].
@@
Recently, Shatashvili et al. [2] studied the nonlinear coupling of circularly polarized EM waves and EAWs in a multi-component relativistic unmagnetized plasma with two groups of electrons (i.e., highly dense degenerate electrons and a small fraction of nondegenerate classical electrons) and immobile positive ions through the description of Zakharov-like equations. They predicted that EM waves could settle into envelope solitons through modulational instability, which can carry essential information about the radiation spectra to be observed far away from the compact objects. The purpose of this work is to extract more findings of the nonlinear coupling between the EM waves and EAWs thereby advancing the existing knowledge of EM solitons by considering a region beyond the plane wave or stable solitonic region in which many linearly modes can be excited as solitary patterns and instead of settling down, saturated to new incoherent patterns after collisions and fusions due to EAW emission. We show that such interactions can undergo through the states of temporal chaos (TC) and spatial partial coherence (SPC), and eventually lead to the emergence of spatiotemporal chaos (STC) [5], which may be an essential signature of EM wave turbulence in the radiational spectra of astrophysical compact objects. The existence of TC and STC and the coexistence of TC and SPC are confirmed by analyzing the Lyapunov exponent spectra [11], the two-point correlation function [12, 13, 9], and the mutual information [14, 15].
@@
The manuscript is organized as follows: In Sec. II, we present the nonlinear Zakharov-like equations, which describe the nonlinear coupling of EM waves and EAWs in multi-component degenerate plasmas and analyze the MI domains and the growth rate in weakly to ultra-relativistic degeneracy regimes. The conditions for the existence of STC are demonstrated in Sec. III. Numerical simulation results of the coupled equations for the evolution of TC, SPC, and STC states are presented in Sec. IV. Finally, Sec. V is left to summarize and conclude the results.
II Basic equations
@@
We consider the nonlinear interaction of hf circularly polarized EM waves and lf EAWs in unmagnetized relativistic plasmas with two groups of electrons, namely relativistic fully degenerate dense electrons and a sparse population of relativistic nondegenerate classical electrons and stationary positive ions. The dynamics of these coupled waves is described by the following set of Zakharov-like equations [2].
@@
Here, $\omega_{\rm{ed}}=\sqrt{4\pi n_{d0}e^{2}/m}$ is the plasma oscillation frequency of degenerate electrons, $\Omega^{2}_{d}=\omega_{\rm{ed}}^{2}/\sqrt{1+R_{0}^{2}}$, $\alpha=n_{\rm{cl0}}/n_{\rm{d0}}$ ($<<1$) is the ratio of the equilibrium number densities of classical and degenerate electrons, $R_{0}\equiv(n_{\rm{d0}}/n_{\rm{cr}})^{1/3}=\hbar(3\pi^{2}n_{\rm{d0}})^{1/3}/mc$ is the electron degeneracy parameter with $n_{\rm{cr}}=m^{3}c^{3}/3\pi^{2}\hbar^{3}\approx 6\times 10^{29}~{}\rm{cm}^{-3}$ denoting the critical number density at which the Fermi momentum, $p_{F}=mc$, i.e., the relativistic degeneracy effect starts. Also, $c_{s}$ is the electron-acoustic speed, given by, $c_{s}^{2}=\alpha c^{2}R_{0}^{2}/3\sqrt{1+R_{0}^{2}}$, $V_{g}~{}(=d\omega_{0}/dk_{0})$ is the group velocity of EM wave envelopes and $b_{1}=(1-\kappa^{2}/\sqrt{R_{0}^{2}+1})$, $b_{2}=\alpha+\kappa^{2}/(R_{0}^{2}+1)^{3/2}$, and $b_{3}=(\sqrt{1+R_{0}^{2}}-\kappa^{2})/R_{0}^{2}$ with $\kappa^{2}=1-R_{0}^{2}/3(1+R_{0}^{2})$ pertain to the coefficients of the nonlocal nonlinear, cubic nonlinear, and ponderomotive nonlinear terms respectively. From the dependency of $c_{s}$ on $\alpha$ (without which $c_{s}$ would vanish or may not be finite) it is evident that a small fraction of nondegenerate electrons is the prerequisite for the generation of low-frequency longitudinal EAWs simultaneously with the EM waves [2]. Equations (1) and (2) were derived using a set of relativistic fluid equations for electrons with the isothermal equation of state for classical electrons and the Fermi-Dirac pressure law for degenerate species and the Maxwell’s equations (See, Ref. [2] for details) to model the dynamics of EM wave radiation ($X$-ray and $\gamma$-ray pulses) emanating from compact astrophysical objects (e.g., white dwarf stars) and interacting with EAWs supported by degenerate dense plasmas surrounding the objects.
@@
where “c.c.” denotes the complex conjugate of the corresponding quantity. Next, we modulate both the amplitude and phase of the EM envelope by the plane wave perturbations as
$\begin{split}&a(x,t)=a_{0}+\tilde{a}\exp(ikx-i\omega t)+\rm{c.c.},\\ &\theta(x,t)=\theta_{0}+\tilde{\theta}\exp(ikx-i\omega t)+\rm{c.c.},\\ \end{split}$ (9)
where $a(x,t)$ and $\theta(x,t)$ are slowly varying functions of $x$ and $t$ such that $\tilde{a}<<a_{0}$, $\tilde{\theta}<<\theta_{0}$. Substituting Eqs. (7) and (8) [using the ansatz (9)] into Eqs. (4) and (5), and looking for nonzero solutions of the amplitudes of perturbations, we obtain the following dispersion relation [2]
@@
Equation (11) for $\Gamma$ agrees with Ref. [2] after one replaces the factors $2$ and $3$ by $3$ and $2$ respectively in the first term under the square brackets. From Fig. 1, it is seen that there are two instability subdomains of $k$: (i) $0<k\lesssim k_{m}$ in which the growth rate $(\Gamma)$ increases with $k$ and reaches a maximum value at $k=k_{m}$ and (ii) $k_{m}<k\lesssim k_{c}$ where the growth rate reduces with $k$ and vanishes at $k=k_{c}$. The latter thus defines a bifurcation point at which the pitchfork bifurcation occurs for unstable $(k<k_{c})$ and stable $(k>k_{c})$ regions. It is also noted that the growth rate of instability and the instability domain are significantly reduced in the regimes of weakly relativistic degeneracy (See the dash-dotted magenta line corresponding to a value of the relativistic degeneracy parameter $R_{0}=0.1$). However, since the degeneracy effect starts playing a role for $R_{0}\gtrsim 1$, i.e., as one enters from moderate to ultra-relativistic degenerate regimes, the instability growth rate tends to decrease, but the instability domain expands with cut-offs at higher values of the wave number of modulation $k$ (See the solid, dashed, and dotted lines). We note that higher values of $R_{0}$ correspond to higher density regimes of degenerate species, and as the value of $R_{0}$ increases, both the group velocity dispersion $(\propto B_{0})$ and the ponderomotive nonlinearity $(\propto b_{3})$ tend to decrease, however, the nonlocal and the cubic nonlinearities get enhanced. As a result, the effective nonlinearity [approximately the cubic nonlinearity proportional to $B_{2}-3B_{1}b_{3}$ for slowly varying envelopes, which is equivalent to taking $\partial^{2}_{t}N\ll\partial^{2}_{x}N$ in Eq. (5)] tends to get reduced, resulting in a reduction of the instability growth rate for the perturbed mode with a cut-off at a higher value of the wave number $k$. Typically, envelope solitons are formed due to MI, and when linearly excited modes at small scales recognize themselves into solitons, the MI gets saturated. In contrast, if the length scale of linearly excited modes is much larger than the most unstable (with maximum growth rate) ones, the EM solitons may be significantly influenced as they interact with the EAWs. Thus, in the high-density regimes $(R_{0}>1)$, there may be the possibilities of collisions and fusions among the modes due to EAW emission [5].
III Conditions for the existence of spatiotemporal chaos

### v2

@@ Abstract
We numerically study the nonlinear interactions of high-frequency circularly polarized electromagnetic (EM) waves and low-frequency electron-acoustic (EA) density perturbations driven by the EM wave ponderomotive force in relativistic plasmas (moderate, strong, and ultra-relativistic) with two groups of electrons–the population of relativistic degenerate dense electrons (bulk plasma) and the sparse relativistic nondegenerate (classical) electrons, and immobile singly charged positive ions. By pattern selection, we show that many solitary patterns can be generated and drenched through modulational instability of EM waves at different spatial length scales and that the EM wave radiation spectra emanating from compact astrophysical objects may not settle into stable envelope solitons but into different incoherent states, including the emergence of temporal and spatiotemporal chaos due to collisions and fusions among the patterns with strong EA wave emission. The appearance of these states is confirmed by analyzing the Lyapunov exponent spectra, correlation function, and mutual information as quantitative evidence. As a result, the redistribution of wave energy from initially exciting many solitary patterns at large scales to a few new incoherent patterns with small wavelengths in the system occurs, leading to the onset of turbulence in astrophysical plasmas.
keywords:
[…]
1 Introduction
@@
The nonthermal radiation of intense electromagnetic (EM) waves (radio wave, $X$-rays, or $\gamma$-rays) can be observed from pulsating stars and during their collapse due to compression of their magnetospheres and a significant increase in the magnetic field. The electric field in this process accelerates charged particles, resulting in charged particle radiation when they move in the magnetic field [kryvdyk2009]. Thus, strong high-frequency (hf) EM waves and high-density relativistic degenerate dense plasmas associated with compact objects (e.g., white dwarfs, neutron stars, active galactic nuclei, etc.) can be simultaneously produced. Such an intense EM wave could play a crucial role in characterizing the nature of the final radiation spectra that will emerge from these compact objects. In the latter, plasmas may consist of high-density relativistic degenerate electrons, contaminated by a small fraction of classical or non-degenerate electrons and positive ions forming only the neutralizing background plasma. They thus can support the propagation of low-frequency (lf) electron-acoustic waves (EAWs) [shatashvili2020nonlinear, misra2021]. Since hf EM fields can scatter into lf EM waves on EAWs, the EM spectral range will change from the original radiation and carry essential information while interacting with EAWs for the self-modulation, soliton formation [shatashvili2020nonlinear, roy2022electromagnetic], and the emergence of spatiotemporal chaos [banerjee2010spatiotemporal], leading to turbulence in astrophysical plasmas. Several authors have studied the nonlinear interactions of hf EM waves with electrostatic lf density perturbations for generating electromagnetic envelope solitons and wakefields [holkundkar2018, roy2020]. However, less effort has been paid to study the emergence of spatiotemporal chaos or EM wave turbulence due to these interactions in high-density regimes.
Typically, the fluid turbulence is associated with chaotic and multi-scale fluid flows. This occurs when fluid kinetic energy becomes high enough to overcome the dissipative effects caused by, for example, fluid viscosity, leading to a transition from large to small-scale eddies through vortex stretching in the energy cascade. On the other hand, in the case of optical turbulence, the key physical mechanisms involve nonlinear self-focusing, which causes the high-intensity light beam to break into filaments, and modulational instability that leads to the formation of envelope solitons in dispersive media. A similar mechanism can also be applicable to plasma media. In planetary ionospheres, plasma turbulence can occur similarly to the fluid turbulence discussed above or due to instabilities in plasmas. However, compared to solar wind or magnetospheric particles, such plasma turbulence may involve wave-particle interactions.
@@
Electromagnetic wave turbulence in astrophysical environments can involve interactions of several physical processes, including shocks, instabilities, and reconnection, and these interactions can result in some observable phenomena like plasma flares and jets, plasma eruptions, and sunspots. The EM wave turbulence can also be generated by other processes like linear conversion of wave modes on plasma density fluctuations, which involve resonant three-wave interactions, and fusion of electrostatic or EM waves. However, several attempts have also been made to understand wave turbulence in terms of nonlinear dynamics and chaos, which involve pattern formation and extensive and incoherent pattern dynamics in both space and time [he2000pattern, he2002harmonic]. We will follow this procedure in the present investigation. In this process, the most significant conservative version of wave turbulence is the Zakharov equations (See, e.g., [banerjee2010spatiotemporal]) that couple hf slowly varying electrostatic (e.g., Langmuir wave) or EM waves (linearly or circularly polarized) to slow electrostatic plasma density perturbations (e.g., electron-acoustic wave and ion-acoustic wave). Typically, decay processes (which may be an induced emission of some waves by others or nonlinear interactions of waves that result in the break-up of waves into others) accumulate wave energy into perturbations with long wavelengths. However, when this wave energy exceeds the modulational instability (MI) threshold, a transfer of energy or a redistribution of energy from initially excited large-scale modes to higher harmonic modes with small wavelengths may be possible. It is believed that some sort of chaotic process in a subsystem may be responsible for driving the energy transfer [banerjee2010spatiotemporal]. Nevertheless, even when the MI of wave envelopes is still present and the energy threshold is exceeded, the energy transfer may not occur, e.g., in the description of wave envelopes by a nonlinear Schröinger (NLS) equation. The latter is known to be an integrable version of the Zakharov equations in the adiabatic limit (very slow-time scales) in which case the MI gets saturated in the formation of envelope solitons [rizzato1998solitons].
Recently, Shatashvili et al. [shatashvili2020nonlinear] studied the nonlinear coupling of circularly polarized EM waves and EAWs in a multi-component relativistic unmagnetized plasma with two groups of electrons (i.e., highly dense degenerate electrons and a small fraction of nondegenerate classical electrons) and immobile positive ions through the description of Zakharov-like equations. They predicted that even with a small fraction of non-degenerate classical electrons, electromagnetic wave envelopes can undergo modulational instability. They also obtained the instability growth rate, which is highly dependent on the degeneracy parameter. Also, they explored the existence of stationary soliton solutions in the adiabatic limit. These solutions can characterize the radiation spectra emanating from compact objects. However, they did not study in detail the influence of the degeneracy parameter $R_{0}$ on the profiles of the instability growth rate or the existence regimes for stable and unstable EM solitons. Their work was limited to the adiabatic limits, specifically in the solitonic region where the length scale of EM wave excitation is assumed to be small and the MI gets saturated. Beyond this scale, solitons may no longer propagate as stable structures due to the strong influence of EAW emission, potentially leading to temporal and spatiotemporal chaos. The existence of these chaotic states, which result from the redistribution of wave energy from large to small-scale perturbations, will carry essential information about the radiation spectra observed far from compact objects. This was not studied before.
@@
The purpose of this work is to extract more findings of the nonlinear coupling between the EM waves and EAWs thereby advancing the existing knowledge of EM solitons as reported in [shatashvili2020nonlinear] by considering a region beyond the plane wave or stable solitonic region in which many linearly modes can be excited as solitary patterns and instead of settling down, saturated to new incoherent patterns after collisions and fusions due to EAW emission. We examine the dynamics across moderate, strong, and ultra-relativistic regimes and show that such interactions can undergo through the states of temporal chaos (TC) and spatial partial coherence (SPC), and eventually lead to the emergence of spatiotemporal chaos (STC) [banerjee2010spatiotemporal]. The latter may be an essential signature of EM wave turbulence in the radiational spectra of astrophysical compact objects. Observing these turbulent signatures is expected to provide insights into the fundamental physics associated with these extreme environments, including particle acceleration and jet formation [Li2025]. Furthermore, the nonlinear and chaotic dynamics will enhance our understanding of the dynamics of degenerate relativistic plasmas and observational signatures, such as spectral broadening, EM pulse fragmentation, and temporal variability, from compact astrophysical objects. The existence of TC and STC and the coexistence of TC and SPC are confirmed by analyzing the Lyapunov exponent spectra [rosenstein1993practical], the two-point correlation function [hohenberg1989chaotic, cai2000chaotic, he2002harmonic], and the mutual information [blahut1987principles, cai2001spatiotemporal], providing a dynamical route to EM wave turbulence with astrophysical implications.
@@
The manuscript is organized as follows: In Sec. 2, we present the nonlinear Zakharov-like equations, which describe the nonlinear coupling of EM waves and EAWs in multi-component degenerate plasmas and analyze the MI domains and the growth rate in weakly to ultra-relativistic degeneracy regimes. The conditions for the existence of STC are demonstrated in Sec. 3. Numerical simulation results of the coupled equations for the evolution of TC, SPC, and STC states are presented in Sec. 4. Finally, Sec. 5 is left to summarize and conclude the results.
2 Basic equations
One-dimensional Zakharov-like equation have been extensively studied in the context of soliton formation and emergence of chaos due to slow response of electrostatic perturbations [see, e.g., Refs. [he2000pattern, he2002harmonic]]. While one-dimensional Zakharov like model has been known to be well-established model for the generation of envelope soliton through modulational instability and chaos or turbulence, two or multi-dimensional model may be useful to study wave collapse or turbulence. Since the modulational instability typically depends on the dispersion and nonlinear coefficients not the spatial coordinates, it is reasonable to consider one space dimension for such instability. The Zakharov-like model can become quasi one-dimensional in those astrophysical conditions where EM wave turbulence is confined to a narrow, elongated structure. Such situations may arise in the environments of compact objects, where the strong magnetic field and field-aligned electron beams restrict the wave dynamics to propagate predominantly along the magnetic field lines, effectively reducing the complex multidimensional wave dynamics to a simpler, quasi-one-dimensional Zakharov-like model.
@@ 2 Basic equations
We aim to consider a one-dimensional model that was previously derived by Shatashvili et al. [shatashvili2020nonlinear] in one-dimensional geometry in the context of MI of circularly polarized EM wave envelopes, and to advance the theory of considering the wave-wave interactions leading to STC. So the possibility of 2D/3D instability of solitary patterns was ruled out. To start with, we consider the nonlinear interaction of hf circularly polarized EM waves and lf EAWs in unmagnetized relativistic plasmas with two groups of electrons, namely relativistic fully degenerate dense electrons and a sparse population of relativistic nondegenerate classical electrons and stationary positive ions. The dynamics of these coupled waves is described by the following set of one-dimensional Zakharov-like equations [shatashvili2020nonlinear].
@@
Here, $\omega_{\rm{ed}}=\sqrt{4\pi n_{d0}e^{2}/m}$ is the plasma oscillation frequency of degenerate electrons, $\Omega^{2}_{d}=\omega_{\rm{ed}}^{2}/\sqrt{1+R_{0}^{2}}$, $\alpha=n_{\rm{cl0}}/n_{\rm{d0}}$ (<<1<<1) is the ratio of the equilibrium number densities of classical and degenerate electrons, $R_{0}\equiv(n_{\rm{d0}}/n_{\rm{cr}})^{1/3}=\hbar(3\pi^{2}n_{\rm{d0}})^{1/3}/mc$ is the electron degeneracy parameter with $n_{\rm{cr}}=m^{3}c^{3}/3\pi^{2}\hbar^{3}\approx 6\times 10^{29}~\rm{cm}^{-3}$ denoting the critical number density at which the Fermi momentum, $p_{F}=mc$, i.e., the relativistic degeneracy effect starts. Also, $c_{s}$ is the electron-acoustic speed, given by, $c_{s}^{2}=\alpha c^{2}R_{0}^{2}/3\sqrt{1+R_{0}^{2}}$, $V_{g}~(=d\omega_{0}/dk_{0})$ is the group velocity of EM wave envelopes and $b_{1}=(1-\kappa^{2}/\sqrt{R_{0}^{2}+1})$, $b_{2}=\alpha+\kappa^{2}/(R_{0}^{2}+1)^{3/2}$, and $b_{3}=(\sqrt{1+R_{0}^{2}}-\kappa^{2})/R_{0}^{2}$ with $\kappa^{2}=1-R_{0}^{2}/3(1+R_{0}^{2})$ pertain to the coefficients of the nonlocal nonlinear, cubic nonlinear, and ponderomotive nonlinear terms respectively. We mention that in Ref. [shatashvili2020nonlinear], the circular polarization of EM waves was considered in the $xy$ plane and the propagation direction along the $z$-axis. However, in the model equations (1) and (2), we have considered the spatial dependence along the $x$-axis for convenience to mean that the circular polarization of EM waves is in the $yz$ plane and the propagation direction along the $x$-axis.
In Eq. (1), $b_{1}$ is the coefficient of nonlocal nonlinearity. Typically, nonlocal effects significantly modify the soliton’s stability and structure in contrast to the cubic or Kerr nonlinearity, which favors the formation of envelope solitons. Such nonlocal effects get enhanced due to an increase in the degeneracy parameter $R_{0}$. This means solitons may not be stable in strong or ultra-relativistic degenerate regimes, but can undergo chaotic states. As the nonlocal nonlinearity enhances and dominates the cubic nonlinearity, the soliton becomes distorted due to collisions and fusions among the patterns by the influence of the electron-acoustic wave emission. This will eventually lead to faster redistribution of wave energy from large to small-scale solitons of higher harmonics.
From the dependency of $c_{s}$ on $\alpha$ (without which $c_{s}$ would vanish or may not be finite) it is evident that a small fraction of nondegenerate electrons is the prerequisite for the generation of low-frequency longitudinal EAWs simultaneously with the EM waves [shatashvili2020nonlinear]. Equations (1) and (2) were derived using a set of relativistic fluid equations for electrons with the isothermal equation of state for classical electrons and the Fermi-Dirac pressure law for degenerate species and the Maxwell’s equations (See, Ref. [shatashvili2020nonlinear] for details) to model the dynamics of EM wave radiation ($X$-ray and $\gamma$-ray pulses) emanating from compact astrophysical objects (e.g., white dwarf stars) and interacting with EAWs supported by degenerate dense plasmas surrounding the objects.
@@
where “c.c." denotes the complex conjugate of the corresponding quantity. Next, we modulate both the amplitude and phase of the EM envelope by the plane wave perturbations as
[…]
where $a(x,t)$ and $\theta(x,t)$ are slowly varying functions of $x$ and $t$ such that $\tilde{a}<<a_{0}$, $\tilde{\theta}<<\theta_{0}$. Substituting Eqs. (7) and (8) [using the ansatz (9)] into Eqs. (4) and (5), and looking for nonzero solutions of the amplitudes of perturbations, we obtain the following dispersion relation [shatashvili2020nonlinear]
@@
Equation (11) for $\Gamma$ agrees with Ref. [shatashvili2020nonlinear] after one replaces the factors $2$ and $3$ by $3$ and $2$ respectively in the first term under the square brackets. From Fig. 1, it is seen that there are two instability subdomains of $k$: (i) $0<k\lesssim k_{m}$ in which the growth rate $(\Gamma)$ increases with $k$ and reaches a maximum value at $k=k_{m}$ and (ii) $k_{m}<k\lesssim k_{c}$ where the growth rate reduces with $k$ and vanishes at $k=k_{c}$. The latter thus defines a bifurcation point at which the pitchfork bifurcation occurs for unstable $(k<k_{c})$ and stable $(k>k_{c})$ regions. It is also noted that the growth rate of instability and the instability domain are significantly reduced in the regimes of weakly relativistic degeneracy (See the dash-dotted magenta line corresponding to a value of the relativistic degeneracy parameter $R_{0}=0.1$). However, since the degeneracy effect starts playing a role for $R_{0}\gtrsim 1$, i.e., as one enters from moderate to ultra-relativistic degenerate regimes, the instability growth rate tends to decrease, but the instability domain expands with cut-offs at higher values of the wave number of modulation $k$ (See the solid, dashed, and dotted lines). We note that higher values of $R_{0}$ correspond to higher density regimes of degenerate species, and as the value of $R_{0}$ increases, both the group velocity dispersion $(\propto B_{0})$ and the ponderomotive nonlinearity $(\propto b_{3})$ tend to decrease, however, the nonlocal and the cubic nonlinearities get enhanced. As a result, the effective nonlinearity [approximately the cubic nonlinearity proportional to $B_{2}-3B_{1}b_{3}$ for slowly varying envelopes, which is equivalent to taking $\partial^{2}_{t}N\ll\partial^{2}_{x}N$ in Eq. (5)] tends to get reduced, resulting in a reduction of the instability growth rate for the perturbed mode with a cut-off at a higher value of the wave number $k$. Typically, envelope solitons are formed due to MI, and when linearly excited modes at small scales recognize themselves into solitons, the MI gets saturated. In contrast, if the length scale of linearly excited modes is much larger than the most unstable (with maximum growth rate) ones, the EM solitons may be significantly influenced as they interact with the EAWs. Thus, in the high-density regimes $(R_{0}>1)$, there may be the possibilities of collisions and fusions among the modes due to EAW emission [banerjee2010spatiotemporal]. From Fig. 1, we also note that as $R_{0}~(>1)$ increases, $k_{m}$ shifts to higher values of $k$, resulting expansions of the instability domains.
3 Conditions for the existence of spatiotemporal chaos